# The large scale geometry of inverse semigroups and their maximal group images

### v1

@@ 1. Introduction
For the distortion to be bounded uniformly across Schützenberger graphs is to say that the natural map from the inverse monoid to its maximal group image is a coarse embedding, in the sense of coarse geometry. This immediately implies that certain geometric properties which are coarse invariant and inherited by subspaces (such as for example Yu’s Property A) can be lifted from the maximal group image to the inverse semigroup. For the special class of $E$-unitary inverse monoids, it was shown in [17] that Property A lifts in this way even without assuming any bound on the distortion. The proof given in [17] was rather indirect, proceeding via $C^{*}$-algebras; having briefly recapped relevant background in Section 2, in Section 3 we digress slightly from our main purpose to answer a question posed in [17] by giving an elementary direct proof of this fact. In fact we establish the stronger statement that Property A lifts through any inverse semigroup morphism which is uniformly bounded-to-one when restricted to $\mathcal{R}$-classes.
@@ 2. Inverse Semigroups and Their Geometry
[…]
@@
Every inverse semigroup has a maximal group image, which can be defined as the quotient by the least congruence which identifies all the idempotents. This congruence is typically denoted by $\sigma$, and we also denote the natural homomorphism from the inverse semigroup to the maximal group image by $\sigma$. Since groups have a well-developed geometric theory, a recurring idea in geometric inverse semigroup theory [12, 17] is to understand when large scale geometric properties of the maximal group image lift to the inverse semigroup. In complete generality, there is not much one can expect – for example if the inverse semigroup has a 00, the maximal group image is trivial. An inverse monoid is called $E$-unitary if the map $\sigma\colon S\to S/\sigma$ is injective on each $\mathrel{\mathcal{R}}$-class. In this case, $\sigma$ induces an embedding of each Schützenberger graph $S\Gamma(s)$ into the Cayley graph of the maximal group image, taken with respect to the same generating set $X$. An equivalent characterization of $E$-unitarity is $\sigma$ being idempotent pure, that is, having the property that $\sigma^{-1}(1)=E(S)$.
@@
We will consider all finitely generated (and more generally all quasi-countable) inverse semigroups to be equipped with a natural extended metric. An extended metric on a set $X$ is a function $d\colon X\times X\to[0,\infty]$ satisfying the usual properties of the metric with respect to the usual addition and partial order on $[0,\infty]$; an extended metric space is a set equipped with an extended metric. We can define an equivalence relation on any extended metric space by putting $x\sim y$ if $d(x,y)<\infty$; the corresponding equivalence classes are called the components of the space. We denote the ball of radius $r$ around the point $x$ by $B(x,r)$. An extended metric space is called uniformly discrete if there is a positive lower bound on the distance between pairs of distinct points, and has bounded geometry if for every $r>0$ there is a finite upper bound on the cardinality of balls of radius $r$.
@@
A map $f:(X,d_{X})\to(Y,d_{Y})$ between (extended) metric spaces is called bornologous if there is a function $\rho_{+}:[0,\infty]\to[0,\infty]$ with $\rho_{+}^{-1}(\infty)=\infty$ such that
@@
for all $x,z\in X$. The map $f$ is a coarse embedding if there is also a function $\rho_{-}:[0,\infty]\to[0,\infty]$ with $\rho_{-}(r)\to\infty$ as $r\to\infty$ such that
@@
[…]
@@
Given a set $X$, an $X$-graph is a directed graph edge-labeled by the set $X\cup X^{-1}$ where each edge labeled by $x\in X$ and has an opposite, inverse edge with label $x^{-1}$, and vice versa. In particular, both group Cayley graphs and inverse semigroup Schützenberger graphs are $X$-graphs, where $X$ is the set of generators.
An $X$-graph is called deterministic if no two edges with the same label have the same initial vertex. Given a not necessarily deterministic, finite $X$-graph, we can always obtain a deterministic quotient by iteratively identifying any pair of edges (as well as their respective inverse pairs) which share the same label and the same initial vertex. We call this process determinization. Schützenberger graphs are of course always deterministic.
Given an inverse monoid $S$ generated by $X$ and a word $w\in(X\cup X^{-1})^{\ast}$ we write $[w]_{S}$ for the element of $S$ defined by $w$. We define the Schützenberger graph of $w$ to be $S\Gamma([w]_{S})$, and sometimes denote it by $S\Gamma(w)$. This is viewed as a birooted graph with marked vertices $\alpha=[ww^{-1}]_{S}$ and $\beta=[w]_{S}$. A key result of Stephen shows that $u$ labels an $\alpha\to\beta$ path in $S\Gamma(w)$ if and only if $[u]_{S}\geq[w]_{S}$.
@@
  • Let $\operatorname{Lin}(w)$ be the linear graph of $w$, i.e. the birooted $X$-graph consisting of a single directed path labeled by $w$ from $\alpha$ to $\beta$. Put $\Gamma_{0}=\operatorname{Lin}(w)$.
  • Perform a $P$-expansion: if there is a path in $\Gamma_{i}$ labeled by one side of a relation such that there is no parallel path in $\Gamma_{i}$ labeled by the other side of that relation, sew on the missing path. More precisely, if $s=t$ is a relation such that there is a path from $u$ to $v$ labeled by $s$ but no path from $u$ to $v$ labeled by $t$, define $\Gamma_{i}^{+}$ by gluing on a disjoint copy of $\operatorname{Lin}(t)$ to $\Gamma_{i}$ from $u$ to $v$.
@@
  • Determinize $\Gamma_{i}^{+}$ to obtain $\Gamma_{i+1}$.
@@
When the inverse monoid is $E$-unitary, and generated by the same set $X$ as its maximal group image, $\sigma$ embeds the Schützenberger graphs of $S$ into the group Cayley graph, which can be conveniently characterized as subgraphs here. The following lemma is straightforward to prove:
@@
An inverse monoid presentation is called special if all relations are of the form $w=1$ for some $w\in(X\cup X^{-1})^{*}$. A special inverse monoid is one that admits a special presentation. Special inverse monoids received considerable attention over recent years [4, 5, 9, 10, 11], particularly the word problem, which is inherently tied to properties of the Schützenberger graphs. In special inverse monoids, $P$-expansions always attach a loop labeled by a relator $w$ around a vertex. It follows that for every $u\in(X\cup X^{-1})^{\ast}$, the Schützenberger graph $S\Gamma(u)$ is the graph obtained from $\operatorname{Lin}(u)$ by sewing on the Schützenberger graph of $1$ at every vertex, and then determinizing.
@@ 3. Property A in $E$-unitary inverse semigroups
In this section we digress slightly from our main theme of studying distortion, in order to give a direct proof of the converse part of [17, Theorem 4.25], that is, that an $E$-unitary inverse semigroup must have Property A if its maximal group image has Property A. We in fact show a more general statement, namely that Property A lifts under any inverse semigroup homomorphism which is uniformly bounded-to-one across all $\mathrel{\mathcal{R}}$-classes.
@@
Our next aim is to establish that Property A lifts under bounded-to-one weakly contractive maps. We use an approach to defining Property A due originally to Higson and Roe for metric spaces (see for example [21]), which was generalized to extended metric spaces in [17].
@@
In particular, applying to the case where $S$ is a finitely quasi-generated $E$-unitary inverse semigroup and $T$ its maximal group image, we obtain a geometric proof of the implication (4) $\implies$ (1) of [17, Theorem 4.25], thereby answering the question the authors pose after the proof. We remark that the authors in [17] study a more restrictive class of inverse semigroups (which they call finitely labelable) than quasi-countable, as [17] predates the introduction of what we call here standard metrics. We state the result in the full generality of quasi-countable inverse semigroups.
@@ 4. Distortion and (weak) $F$-Inverse properties
Let $S$ be a quasi-countable inverse semigroup and $G$ its maximal group image, both equipped with proper length functions. We say that $S$ has locally bounded group distortion if for every $\mathcal{R}$-class there exists a weakly increasing function $\phi:[0,\infty)\to[0,\infty)$ such that
$l_{S}(s)\leq\phi(l_{G}(s\sigma))\textrm{ for all }s\in S.$
We say that $S$ has uniformly bounded group distortion if it has locally bounded group distortion using the same function $\phi$ for every $\mathcal{R}$-class.
We remark that in the case that $S$ is finitely generated and equipped with a word metric, it is equivalent to consider in the definition functions $\phi:\mathbb{N}\to\mathbb{N}$. We also note that one may, in the definition, require without loss of generality that the function $\phi$ be increasing. Indeed, if there is a $\phi$ satisfying the given inequalities then clearly we may assume $\phi$ is supported only on the image of $l_{G}$. Now since $l_{G}$ is assumed to be proper its image intersects each bounded set in finitely many points, so $\phi$ is supported on only finitely many points in each bounded set. It follows easily that $\phi$ is bounded above by some increasing function, which we may use to replace it.
Recall that here, as usual, we consider both $S$ and its maximal group image equipped with standard metrics. The following proposition will allow us (among other things) to conclude that the definitions above are independent of the particular standard metrics chosen. The reader may prefer always to think of the metrics being weighted path metrics with respect to countable generating sets, or even to restrict attention to the finitely generated case and assume the metric to be the true path metric in the Schützenberger graphs.
@@
Assume first that $S$ has uniformly bounded group distortion, with $\phi$ being the bounding function. By the remark under Definition 4.1 we may assume that $\phi$ is increasing. By Proposition 3.1 the map is bornologous and so there exists a function $\rho_{+}\colon[0,\infty]\to[0,\infty]$ with $\rho_{+}^{-1}(\infty)=\infty$ such that $d_{G}(x\sigma,y\sigma)\leq\rho_{+}(d_{S}(x,y))$ for any $x,y\in S$, and in particular for any pair of $\mathrel{\mathcal{R}}$-related elements $x,y\in S$. What remains is to find a function $\rho_{-}:[0,\infty]\to[0,\infty]$ which tends to infinity and such $\rho_{-}(d_{S}(x,y))\leq d_{G}(x\sigma,y\sigma)$ for every pair of $\mathrel{\mathcal{R}}$-related elements $x,y\in S$. We define $\rho_{-}$ by
@@
It is easy to see that $\rho_{-}$ is weakly increasing, and since $\phi$ is weakly increasing by assumption, we must have $\rho_{-}(k)\geq p$ whenever $k>\phi(p)$, so $\rho_{-}$ tends to infinity. Now for every $r\in\mathbb{R}$ we have
@@
since the set $\{p\mid\phi(r)\leq\phi(p)\}$ contains $r$. For every $s\in S$, by the defining property of $\phi$ we have $l_{S}(s)\leq\phi(l_{G}(s\sigma))$, so applying the weakly increasing function $\rho_{-}$ to both sides we obtain
$\rho_{-}(l_{S}(s))\leq\rho_{-}(\phi(l_{G}(s\sigma)))\leq l_{G}(s\sigma).$
So for every pair $x,y$ of $\mathrel{\mathcal{R}}$-related elements,
$\rho_{-}(d_{S}(x,y))=\rho_{-}(l_{S}(x^{-1}y))\leq l_{G}((x^{-1}y)\sigma)=d_{G}(x\sigma,y\sigma)$
@@
Conversely, assume that $\sigma$ is a coarse embedding on each $\mathrel{\mathcal{R}}$-class with respect to uniformly chosen functions $\rho_{-}$ and $\rho_{+}$, that is we have $\rho_{-}(d_{S}(x,y))\leq d_{G}(x\sigma,y\sigma)$ for any pair of $\mathrel{\mathcal{R}}$-related elements $x,y\in S$. Then for every $s\in S$, $\rho_{-}(l_{S}(s))=\rho_{-}(d_{S}(s,ss^{-1}))\leq d_{G}(s\sigma,1)=l_{G}(s\sigma)$. Now since $\rho_{-}$ tends to infinity we may define $\phi$ by
@@
whereupon $\phi(\rho_{-}(k))\geq k$ for all $k\in\mathbb{R}$. Moreover, $\phi$ is easily seen to be weakly increasing, so for every $s$ we have
$l_{S}(s)\leq\phi(\rho_{-}(l_{S}(s))\leq\phi(l_{G}(s\sigma))$
@@
We remark that there is a subtle, but significant difference between $\sigma$ being a coarse equivalence (which implies that $S$ is a group) and $S$ having uniformly bounded group distortion, the former corresponding to an inequality $\rho_{-}(d_{S}(x,y))\leq d_{G}(x\sigma,y\sigma)\leq\rho_{+}(d_{S}(x,y))$ for all pairs of $x,y$, and the latter only satisfying this when $x\mathrel{\mathcal{R}}y$.
@@
The definitions of locally and uniformly bounded group distortion are independent of the choice of standard metric.
The fact that a quasi-countable inverse semigroup can fail to have locally bounded group distortion was observed (phrased in the language of course embeddings) in [17], with an example given in [20, Proposition 5.4.6]. Indeed, the following construction will allow us to show something even stronger: that every subgraph of a group Cayley graph is the image of such a map, for a suitable Schützenberger graph of a suitable $E$-unitary inverse monoid. (This is already known for finite subgraphs as a consequence of the Margolis-Meakin expansion construction [18], but for infinite subgraphs it does not seem to have been previously observed. In the related statement [2, Theorem 3.24], it is observed that every uniformly discrete, bounded metric space arises, up to bijective coarse equivalence, as the $\mathrel{\mathcal{R}}$-class of some inverse semigroup, but the construction yields a non-$E$-unitary inverse semigroup.)
@@
    $S$ has a Schützenberger graph which is isomorphic to $\Delta$ with an added $e$-edge labeling a loop around a chosen vertex.
  • Proof.
@@
      $w^{2}=w$ for every word $w\in(X\cup X^{-1})^{\ast}$ with $[w]_{G}=1$;

### v2

@@ 1. Introduction
If the distortion is uniformly bounded across all Schützenberger graphs, then certain geometric properties which are coarse invariant and inherited by subspaces (such as for example Yu’s Property A) can be lifted from the maximal group image to the inverse semigroup. For the special class of $E$-unitary inverse monoids, it was shown in [16] that Property A lifts in this way even without assuming any bound on the distortion. The proof given in [16] was rather indirect, proceeding via $C^{*}$-algebras; having briefly recapped relevant background in Section 2, in Section 3 we digress slightly from our main purpose to answer a question posed in [16] by giving an elementary direct proof of this fact. In fact we establish the stronger statement that Property A lifts through any inverse semigroup morphism which is uniformly bounded-to-one when restricted to $\mathcal{R}$-classes.
@@ 2. Inverse Semigroups and Their Geometry
Given an inverse semigroup $S$ generated by $X$, we define the Cayley graph of $S$ as usual: the vertex set is $S$, and the edges are of the form $s\xrightarrow{x}sx$ for $s\in S,x\in X\cup X^{-1}$. The strongly connected component of the Cayley graph containing the vertex $s$ is called the Schützenberger graph of $s$, and is denoted by $S\Gamma(s)$. We define an equivalence relation on $S$ by $s\mathrel{\mathcal{R}}t$ if and only if $s$ and $t$ lie in the same Schützenberger graph, or equivalently, if and only if $sS=tS$. (The relation $\mathrel{\mathcal{R}}$ is one of the five equivalence relations defined by Green which play a central role most aspects of semigroup theory [13], but in this article we shall have no need for the other relations.) It is an important fact that $sx\mathrel{\mathcal{R}}s$ implies $sxx^{-1}=s$, which means that the edges of Schützenberger graphs occur in inverse pairs: there is an edge from $u$ to $v$ labelled $x$ if and only if there is an edge from $v$ to $u$ labelled $x^{-1}$. It follows that the path metric on these graphs defines a metric on each $\mathrel{\mathcal{R}}$-class. Another useful observation is that $s\mathrel{\mathcal{R}}ss^{-1}$ for every $s\in S$. Moreover, the $\mathrel{\mathcal{R}}$-relation can be characterised by $s\mathrel{\mathcal{R}}t$ if and only if $ss^{-1}=tt^{-1}$. It follows that each $\mathrel{\mathcal{R}}$-class has a unique idempotent.
@@
Every inverse semigroup has a maximal group image, which can be defined as the quotient by the least congruence which identifies all the idempotents. This congruence is typically denoted by $\sigma$, and we also denote the natural homomorphism from the inverse semigroup to the maximal group image by $\sigma$. Since groups have a well-developed geometric theory, a recurring idea in geometric inverse semigroup theory [12, 16] is to understand when large scale geometric properties of the maximal group image lift to the inverse semigroup. In complete generality, there is not much one can expect – for example if the inverse semigroup has a 0, the maximal group image is trivial. An inverse monoid is called $E$-unitary if the map $\sigma\colon S\to S/\sigma$ is injective on each $\mathrel{\mathcal{R}}$-class. In this case, $\sigma$ induces an embedding of each Schützenberger graph $S\Gamma(s)$ into the Cayley graph of the maximal group image, taken with respect to the same generating set $X$. An equivalent characterisation of being $E$-unitary is $\sigma$ being idempotent pure, that is, having the property that $\sigma^{-1}(1)=E(S)$.
We briefly recall some other definitions from inverse semigroup theory which we shall need later. An inverse semigroup is called $F$-inverse if every $\sigma$-class contains a maximum element under the natural partial order; in fact it can easily be shown that $F$-inverse semigroups are necessarily $E$-unitary inverse monoids. A map $\tau:S\to T$ between inverse monoids $S$ and $T$ is called a premorphism if $\tau(1_{S})=1_{T}$, and for all $s,t\in S$ we have $\tau(s^{-1})=\tau(s)^{-1}$ and $\tau(s)\tau(t)\leq\tau(st)$, where $\leq$ denotes the natural partial order on $T$.
@@
We will consider all finitely generated (and more generally all quasi-countable) inverse semigroups to be equipped with a natural extended metric. An extended metric on a set $X$ is a function $d\colon X\times X\to[0,\infty]$ satisfying the usual properties of the metric with respect to the usual addition and partial order on $0,\infty]$; an extended metric space is a set equipped with an extended metric. We can define an equivalence relation on any extended metric space by putting $x\sim y$ if $d(x,y)<\infty$; the corresponding equivalence classes are called the components of the space. We denote the ball of radius $r$ around the point $x$ by $B(x,r)$. An extended metric space is called uniformly discrete if there is a positive lower bound on the distance between pairs of distinct points, and has bounded geometry if for every $r>0$ there is a finite upper bound on the cardinality of balls of radius $r$.
@@
A map $f:(X,d_{X})\to(Y,d_{Y})$ between (extended) metric spaces is called bornologous if there is a monotonically non-decreasing function $\rho_{+}:[0,\infty]\to[0,\infty]$ with $\rho_{+}^{-1}(\infty)=\infty$ such that
@@
for all $x,z\in X$. The map $f$ is a coarse embedding if there is also a monotonically non-decreasing function $\rho_{-}:[0,\infty]\to[0,\infty]$ with $\rho_{-}(r)\to\infty$ as $r\to\infty$ such that
@@
A canonical example of a standard metric is the word metric on a finitely (quasi-)generated inverse semigroup. More precisely, if $S=\langle X\cup E(S)\rangle$ where $X$ is a finite set, then one can define the distance between two $\mathrel{\mathcal{R}}$-related elements $s$ and $t$ to be their distance in their Schützenberger graph with respect to the path metric, while the distance between elements which are not $\mathrel{\mathcal{R}}$-related is defined to be $\infty$. More generally, if $X$ is countable, one can instead consider a weighted word metric, corresponding to a function $w\colon X\to\mathbb{N}$ where $w^{-1}(n)$ is finite for all $n\in\mathbb{N}$. In this setting, instead of simply taking the path metric in the Schützenberger graphs (which would not result in a proper metric), we take the length of the edges $s\xrightarrow{x}sx$ and $sx\xrightarrow{x^{-1}}s$ to be $w(x)$. More precisely, for each pair of $\mathrel{\mathcal{R}}$-related elements $s,t\in S$, we obtain the distance
@@
Given a set $X$, an $X$-graph is a directed graph edge-labelled by the set $X\cup X^{-1}$ where each edge labelled by $x\in X$ and has an opposite, inverse edge with label $x^{-1}$, and vice versa. In particular, both group Cayley graphs and inverse semigroup Schützenberger graphs are $X$-graphs, where $X$ is the set of generators.
An $X$-graph is called deterministic if no two edges with the same label have the same initial vertex. Given a not necessarily deterministic, finite $X$-graph, we can always obtain a deterministic quotient by iteratively identifying any pair of edges (as well as their respective inverse pairs) which share the same label and the same initial vertex. We call this process determinisation. Schützenberger graphs are of course always deterministic.
Given an inverse monoid $S$ generated by $X$ and a word $w\in(X\cup X^{-1})^{\ast}$ we write $w]_{S}$ for the element of $S$ defined by $w$. We define the Schützenberger graph of $w$ to be $S\Gamma([w]_{S})$, and sometimes denote it by $S\Gamma(w)$. This is viewed as a birooted graph with marked vertices $\alpha=[ww^{-1}]_{S}$ and $\beta=[w]_{S}$. A key result of Stephen shows that $u$ labels an $\alpha\to\beta$ path in $S\Gamma(w)$ if and only if $u]_{S}\geq[w]_{S}$.
@@
  • Let $\operatorname{Lin}(w)$ be the linear graph of $w$, i.e. the birooted $X$-graph consisting of a single directed path labelled by $w$ from $\alpha$ to $\beta$. Put $\Gamma_{0}=\operatorname{Lin}(w)$.
  • Perform a $P$-expansion: if there is a path in $\Gamma_{i}$ labelled by one side of a relation such that there is no parallel path in $\Gamma_{i}$ labelled by the other side of that relation, sew on the missing path. More precisely, if $s=t$ is a relation such that there is a path from $u$ to $v$ labelled by $s$ but no path from $u$ to $v$ labelled by $t$, define $\Gamma_{i}^{+}$ by gluing on a disjoint copy of $\operatorname{Lin}(t)$ to $\Gamma_{i}$ from $u$ to $v$.
@@
  • Determinise $\Gamma_{i}^{+}$ to obtain $\Gamma_{i+1}$.
@@
When the inverse monoid is $E$-unitary, and generated by the same set $X$ as its maximal group image, $\sigma$ embeds the Schützenberger graphs of $S$ into the group Cayley graph, which can be conveniently characterised as subgraphs here. The following lemma is straightforward to prove:
@@
An inverse monoid presentation is called special if all relations are of the form $w=1$ for some $w\in(X\cup X^{-1})^{*}$. A special inverse monoid is one that admits a special presentation. Special inverse monoids received considerable attention over recent years [4, 5, 9, 10, 11], particularly the word problem, which is inherently tied to properties of the Schützenberger graphs. In special inverse monoids, $P$-expansions always attach a loop labelled by a relator $w$ around a vertex. It follows that for every $u\in(X\cup X^{-1})^{\ast}$, the Schützenberger graph $S\Gamma(u)$ is the graph obtained from $\operatorname{Lin}(u)$ by sewing on the Schützenberger graph of $1$ at every vertex, and then determinizing.
@@ 3. Property A in $E$-unitary inverse semigroups
In this section we digress slightly from our main theme of studying distortion, in order to give a direct proof of the converse part of [16, Theorem 4.25], that is, that an $E$-unitary inverse semigroup must have Property A if its maximal group image has Property A. We in fact show a more general statement, namely that Property A lifts under any inverse semigroup homomorphism which is uniformly bounded-to-one across all $\mathrel{\mathcal{R}}$-classes.
@@
Our next aim is to establish that Property A lifts under bounded-to-one weakly contractive maps. We use an approach to defining Property A due originally to Higson and Roe for metric spaces (see for example [20]), which was generalised to extended metric spaces in [16].
@@
In particular, applying to the case where $S$ is a finitely quasi-generated $E$-unitary inverse semigroup and $T$ its maximal group image, we obtain a geometric proof of the implication (4) $\implies$ (1) of [16, Theorem 4.25], thereby answering the question the authors pose after the proof. We remark that the authors in [16] study a more restrictive class of inverse semigroups (which they call finitely labelable) than quasi-countable, as [16] predates the introduction of what we call here standard metrics. We state the result in the full generality of quasi-countable inverse semigroups.
@@ 4. Distortion and (weak) $F$-Inverse properties
Let $S$ be a quasi-countable inverse semigroup and $G$ its maximal group image, equipped with standard metrics $d_{S}$ and $d_{G}$ respectively. We say that $S$ has locally bounded group distortion if for every $\mathcal{R}$-class $R$ there exists a function $\phi:[0,\infty)\to[0,\infty)$ such that
$d_{S}(s,t)\leq\phi(d_{G}(s\sigma,t\sigma))\textrm{ for all }s,t\in R.$
We say that $S$ has uniformly bounded group distortion if it has locally bounded group distortion with the same choice of $\phi$ for every $\mathcal{R}$-class.
We remark that in the case that $S$ is finitely generated and equipped with a word metric, it is equivalent to consider in the definition functions $\phi:\mathbb{N}\to\mathbb{N}$. We also note that one may, in the definition, require without loss of generality that the function $\phi$ be increasing. Indeed, if there is a $\phi$ satisfying the given inequalities then clearly we may assume $\phi$ is supported only on the image of $d_{G}$. Now since $d_{G}$ is assumed to be proper its image intersects each bounded set in finitely many points, so $\phi$ is supported on only finitely many points in each bounded set. It follows easily that $\phi$ is bounded above by some increasing function, which we may use to replace it.
The following proposition will allow us (among other things) to conclude that the definitions above are independent of the particular standard metrics chosen. The reader may prefer always to think of the metrics being weighted path metrics with respect to countable generating sets, or even to restrict attention to the finitely generated case and assume the metric to be the true path metric in the Schützenberger graphs.
@@
Assume first that $S$ has uniformly bounded group distortion, with $\phi$ being the bounding function. By the remark under Definition 4.1 we may assume that $\phi$ is increasing. By Proposition 3.1 the map $\sigma$ is bornologous and so there exists a function $\rho_{+}\colon[0,\infty]\to[0,\infty]$ with $\rho_{+}^{-1}(\infty)=\infty$ such that $d_{G}(s\sigma,t\sigma)\leq\rho_{+}(d_{S}(s,t))$ for any $s,t\in S$, and in particular for any pair of $\mathrel{\mathcal{R}}$-related elements $s,t\in S$. What remains is to find a function $\rho_{-}:[0,\infty]\to[0,\infty]$ which tends to infinity and such $\rho_{-}(d_{S}(s,t))\leq d_{G}(s\sigma,t\sigma)$ for every pair of $\mathrel{\mathcal{R}}$-related elements $s,t\in S$. We define $\rho_{-}$ by
@@
It is easy to see that $\rho_{-}$ is weakly increasing, and since $\phi$ is increasing by assumption, we must have $\rho_{-}(k)\geq p$ whenever $k>\phi(p)$, so $\rho_{-}$ tends to infinity. Now for every $r\in\mathbb{R}$ we have
@@
since the set $\{p\mid\phi(r)\leq\phi(p)\}$ contains $r$. For every $s,t\in S$ with $s\mathrel{\mathcal{R}}t$, by the defining property of $\phi$ we have $d_{S}(s,t)\leq\phi(d_{G}(s\sigma,t\sigma))$, so applying the weakly increasing function $\rho_{-}$ to both sides we obtain
$\rho_{-}(d_{S}(s,t))\leq\rho_{-}(\phi(d_{G}(s\sigma,t\sigma)))\leq d_{G}(s\sigma,t\sigma)$
@@
Conversely, assume that $\sigma$ is a coarse embedding on each $\mathrel{\mathcal{R}}$-class with respect to uniformly chosen functions $\rho_{-}$ and $\rho_{+}$, that is we have $\rho_{-}(d_{S}(s,t))\leq d_{G}(s\sigma,t\sigma)$ for any pair of $\mathrel{\mathcal{R}}$-related elements $s,t\in S$. Now since $\rho_{-}$ tends to infinity we may define $\phi$ by
@@
whereupon $\phi(\rho_{-}(k))\geq k$ for all $k\in\mathbb{R}$. Moreover, $\phi$ is easily seen to be weakly increasing, so for every $s,t\in S$ with $s\mathrel{\mathcal{R}}t$ we have
$d_{S}(s,t)\leq\phi(\rho_{-}(d_{S}(s,t))\leq\phi(d_{G}(s\sigma,t\sigma))$
@@
We remark that there is a subtle, but significant difference between $\sigma$ being a coarse equivalence (which implies that $S$ is a group) and $S$ having uniformly bounded group distortion, the former corresponding to an inequality $\rho_{-}(d_{S}(s,t))\leq d_{G}(s\sigma,t\sigma)\leq\rho_{+}(d_{S}(s,t))$ for all pairs of $s,t$, and the latter only satisfying this when $s\mathrel{\mathcal{R}}t$.
@@
The definitions of locally and uniformly bounded group distortion are independent of the choice of standard metrics.
We can also characterise bounded group distortion in terms of proper length functions, which is the formulation we most often use:
Proposition 4.4.
Let $S$ be a quasi-countable inverse semigroup and $G$ its maximal group image, equipped with proper length functions $l_{S}$ and $l_{G}$ respectively. Then $S$ has locally bounded group distortion if and only if for every $\mathcal{D}$-class $D$ of $S$ there exists a function $\phi:[0,\infty)\to[0,\infty)$ such that for every $s\in D$ we have $l_{S}(s)\leq\phi(l_{G}(s\sigma))$. $S$ has uniformly bounded group distortion if it satisfies the above condition uniformly with the same function $\phi$ for every $\mathcal{D}$-class.
Proof.
Let $d_{S}$ and $d_{G}$ be the standard metrics corresponding to $l_{S}$ and $l_{G}$ as discussed below Definition 2.2. Suppose $S$ has locally bounded group distortion. Fix a $\mathcal{D}$-class $D$ and choose an $\mathcal{R}$-class $R$ in $D$, and a function $\phi:[0,\infty)\to[0,\infty)$ such that $d_{S}(x,y)\leq\phi(d_{G}(x\sigma,y\sigma))$ for all $x,y\in R$. Let $s\in D$ and denote the $\mathrel{\mathcal{R}}$-class of $s$ by $R_{s}$; we aim to show that $l_{S}(s)\leq\phi(l_{G}(s\sigma))$. Then by Green’s Lemma [3, Theorem 2.3], we may choose $p,q\in S$ such that the left translations
$\begin{array}[]{cc}\lambda_{p}\colon R_{s}\to R&x\mapsto px\\ \lambda_{q}\colon R\to R_{s}&y\mapsto qy\end{array}$
are mutually inverse bijections between $R_{s}$ and $R$. Because $d_{S}$ is left subinvariant, these bijections must both be weakly contracting, which since they are mutually inverse implies that both are distance-preserving. Since $ps$ and $pss^{-1}$ lie in $R$, we have
$l_{S}(s)=d_{S}(ss^{-1},s)=d_{S}(pss^{-1},ps)\leq\phi(d_{G}((pss^{-1})\sigma,(ps)\sigma))=\phi(d_{G}(1,s))=\phi(l_{G}(s)),$
where the penultimate equality holds because the metric on $G$ is left invariant.
Conversely, suppose for a $\mathcal{D}$-class $D$ there is a function $\phi:[0,\infty)\to[0,\infty)$ such that $l_{S}(s)\leq\phi(l_{G}(s\sigma))\textrm{ for all }s\in D.$ Then for any $x,y\in D$ with $x\mathrel{\mathcal{R}}y$, we have that $x^{-1}y\in D$, and
$d_{S}(x,y)=l_{S}(x^{-1}y)\leq\phi(l_{G}((x^{-1}y)\sigma))=\phi(d_{G}(x,y)),$
hence $S$ has locally bounded group distortion. The uniformly bounded case follows immediately. ∎
The fact that a quasi-countable inverse semigroup can fail to have locally bounded group distortion was observed (phrased in the language of course embeddings) in [16], with an example given in [19, Proposition 5.4.6]. Indeed, the following construction will allow us to show something even stronger: that every subgraph of a group Cayley graph is the image in the maximal group image of a suitable Schützenberger graph of a suitable $E$-unitary inverse monoid. (This is already known for finite subgraphs as a consequence of the Margolis-Meakin expansion construction [17], but for infinite subgraphs it does not seem to have been previously observed. In the related statement [2, Theorem 3.24], it is observed that every uniformly discrete, bounded metric space arises, up to bijective coarse equivalence, as the $\mathrel{\mathcal{R}}$-class of some inverse semigroup, but the construction yields a non-$E$-unitary inverse semigroup.)
@@
    $S$ has a Schützenberger graph which is isomorphic to $\Delta$ with an added $e$-edge labelling a loop around a chosen vertex.
@@
    $w^{2}=w$ for every word $w\in(X\cup X^{-1})^{\ast}$ with $w]_{G}=1$;